Write hamiltonian for free particle in one dimensional.
1
Expert's answer
2020-10-25T18:31:41-0400

In general case,

"H = p \dot{q} - L"

where p is generalized momentum and "\dot{q}" is generalized velocity and "L" is Lagrange function or Lagrangian. In classical mechanics, Lagrange function "L = T - U" , where T is kinetic energy and U is potential energy.

In case of free particle, there is no potential field, so "\displaystyle L= T = \frac{mv^2}{2} = \frac{m \dot{x}^2}{2}"

"p = mv = m \dot{x}"

"\dot{q} = \dot{x}"

"\displaystyle H = m \dot{x}^2 - \frac{m \dot{x}^2}{2} = \frac{m \dot{x}^2}{2}"

As we see here, Hamiltonian for free particle is coincident with its kinetic energy.